OpenStudy (anonymous):

HEY CAN ANY ONE OUT HERE HELP ME SOLVING THIS PROBLEM LIM(X-0) (SINX) POWER SINX

OpenStudy (espex):

Assuming your equation reads like this: \[\lim_{x \rightarrow 0}\sin (x)^{\sin (x)}\] Given the continuity of sin(x) at x=0 you can rewrite the equation like:\[\sin(x)^{\lim_{x \rightarrow 0}\sin(x)}\] Then evaluate at the limit to get: \[\sin(x)^{\sin(0)} \rightarrow \sin(x)^0 = 1\]